# Sliding Puzzles Gym: A Scalable Benchmark for State Representation in Visual Reinforcement Learning

### v4

@@ 4 Methods
In this section, we detail the evaluation protocol for various RL agents within the SPGym framework, focusing on their ability to manage increasing visual diversity.
@@
We sample images from ImageNet-1k’s validation split (Russakovsky et al., 2015) to construct visual observations, resizing each image to $84\times 84$ pixels and normalizing values to $[0,1]$. To isolate visual diversity, we fix the puzzle size to $3\times 3$ and vary only the image pool size $p$. For each training run, we randomly sample $p$ distinct images to create a fixed image pool. At episode start, we randomly select one image from this pool to generate the puzzle observations. We cap the number of environment steps to 10M and limit episodes to 1,000 steps.
@@
For SAC, we begin with the standard implementation for discrete action spaces proposed by Christodoulou (2019). We then examine several representation learning variants: RAD (Laskin et al., 2020a), which employs data augmentation; CURL (Laskin et al., 2020b), which uses contrastive learning; SPR (Schwarzer et al., 2020), which incorporates self-supervised prediction; DBC (Zhang et al., 2021), which focuses on state metric learning; SAC-AE and SAC-VAE (Yarats et al., 2021b), which utilize reconstruction-based learning; and Simple Baseline (SB) (Tomar et al., 2023), which implements a simplified approach with reward and transition prediction.
@@
For SAC-based agents, we follow Yarats et al. (2021b) and Tomar et al. (2023) by blocking actor gradients through the encoder while allowing critic and auxiliary gradients to prevent representation collapse. PPO and DreamerV3 maintain their original gradient flow patterns. While PPO and DreamerV3 worked robustly with default configurations, SAC-based agents required tuning of the temperature parameter $\alpha$ (Appendix A.4.1). We found a fixed value of $0.05$ to work best across all SAC variants, as the automatic tuning from Haarnoja et al. (2018b) proved ineffective here.
Where applicable, we preserve uniform hyperparameter while drawing algorithm-specific configurations from their respective source papers. The publicly available code repository [^2] Accessible at https://github.com/bryanoliveira/spgym-experiments extends both CleanRL (Huang et al., 2022) and the official DreamerV3 codebase. Additional methodological details are provided in Appendix A.
@@ 5 Results
5.1 Can SPGym distinguish agents on representation learning capabilities?
Table 1: Million steps to reach 80% success rate across pool sizes. Lower is better. Best performing variant for each algorithm and pool size is highlighted in bold.
@@ 5.1 Can SPGym distinguish agents on representation learning capabilities?
Table 1 demonstrates SPGym’s ability to differentiate agents based on their representation learning approaches across varying levels of visual diversity. Our primary goal in this analysis is to compare the sample efficiency with which different agents learn useful visual representations. Notably, certain representation learning methods provide clear benefits, though the extent can depend on the pool size.
@@
To explore the potential benefits of leveraging pretrained encoders in this setup, we evaluated PPO with two types of pretrained encoders. It is important to acknowledge that direct comparisons of sample efficiency with encoders trained from scratch can be nuanced, as pretrained encoders have inherently been exposed to significantly more data during their pretraining phase. The in-distribution pretrained encoder (PT (ID)) for PPO, intended as a proxy to estimate an approximate upper bound on performance achievable with pretraining, significantly boosts sample efficiency across all tested pool sizes, almost matching the performance of PPO with one-hot observations (see Appendix C.1.1). The out-of-distribution pretrained encoder (PT (OOD)), designed to represent the average performance one might expect from a generally pretrained encoder, offers more modest gains which also decrease with larger pools.
For SAC, data augmentation via RAD consistently and substantially improves efficiency across all tested pool sizes, demonstrating increasing benefits as visual diversity scales. The improvements are particularly pronounced for larger pools, making SAC significantly faster at handling visual diversity. Conversely, many sophisticated auxiliary representation learning methods struggle in SPGym. For instance, CURL, SPR, and VAE variants consistently require more samples than standard SAC across all tested pool sizes, with particularly poor performance on larger pools. DBC also generally underperforms or offers marginal improvements compared to standard SAC.
@@
DreamerV3 demonstrates particularly strong performance, consistently outperforming both PPO and SAC variants across all pool sizes. The full DreamerV3 model shows remarkably stable performance across different pool sizes, significantly better than most other methods. The variant without decoder gradients shows reduced performance across all configurations, highlighting the importance of the world model’s discrete reconstruction objective. This suggests that DreamerV3’s approach of learning a predictive model of the environment provides a particularly effective foundation for handling visual diversity in SPGym.
@@
These results highlight SPGym’s value as a diagnostic tool, effectively distinguishing which approaches can handle its distinct combination of visual diversity and structured dynamics. The benchmark exposes important limitations in methods that have proven successful in other visual RL domains, suggesting that their underlying assumptions may not align with SPGym’s unique characteristics. For detailed performance analysis and hypothesized explanations for agents behavior, see Appendix C.2.1.
@@ 5.2 How does visual diversity affect agents?
These distinct degradation patterns suggest differences in handling increasing visual diversity, potentially due to agents memorizing features instead of learning generalizable representations, thus exhausting network capacity. PPO’s early failure indicates difficulty forming robust representations and rapid capacity exhaustion. SAC’s better performance on medium pools, possibly aided by its replay buffer and encoder training independent of actor gradients, suggests stronger initial representation learning. However, it also faces fundamental limitations, potentially as memorization becomes unsustainable. DreamerV3’s world model architecture appears to foster more compressed representations, leading to more graceful degradation, but it too struggles with very large pools.
Preliminary DreamerV3 experiments with extremely large image pools (10,000-20,000 images) support this interpretation: agents failed to learn useful policies as each observation became virtually unique. This suggests that increased dataset scale alone is insufficient if the RL signal is too sparse for encoder training amidst high diversity. These methods appear to memorize rather than generalize, struggling when network capacity is overwhelmed by unique inputs. This highlights fundamental limitations of current methods when facing diverse visuals where assumptions break and the learning signal is too weak for generalization.
@@ 5.3 Can agents generalize to unseen visual inputs?
Our analysis of out-of-distribution (OOD) generalization capabilities on PPO and SAC agents reveals a stark contrast between in-distribution and generalization performance. We evaluate two levels of generalization difficulty: ‘Easy’ OOD, using augmented versions of training images, and ‘Hard’ OOD, using completely unseen images. Since meaningful generalization evaluation requires agents that first achieve competent in-distribution performance, we primarily focus our OOD analysis on smaller pool sizes (i.e., 1, 5, and 10) where most methods successfully learn the task.
Table 2: Success rate of PPO and SAC agents on Easy OOD across different pool sizes. Higher is better.
@@
Despite most agents achieving high success rates on their training distribution, they exhibit varying degrees of failure on both generalization challenges. Table 2 illustrates this for the Easy OOD setting across pool sizes 1, 5, and 10, showing a general decline in OOD performance as training pool diversity increases. For Easy OOD evaluation, we applied the augmentations described in Appendix A.3 to training images and tested PPO and SAC agents across 100 episodes for each augmentation type. We observe that while some SAC variants demonstrate improved robustness to these augmentations compared to base SAC, particularly on pool size 1, this Easy OOD performance generally degrades as the training pool size increases. We found a strong correlation (Pearson r=-0.81, p=2.5e-12) between success rates on these augmented images (Easy OOD) and agents’ in-distribution sample efficiency, suggesting that representations robust to simple transformations are tied to faster learning on SPGym.
@@
For Hard OOD evaluation, we test PPO, SAC and Dreamer agents on 100 episodes using images independently sampled from the ImageNet-1k validation split, repeating across 5 seeds. Here, agents almost universally fail, achieving near 0% success rates across all methods and training configurations. This stark failure on completely unseen images is a crucial diagnostic finding, exposing fundamental limitations of current end-to-end RL methods when faced with visual inputs that differ substantially from their training distribution. The complete inability to generalize suggests that what agents learn to achieve good task performance is primarily memorization of specific visual patterns rather than learning generalizable representations.
@@
Intuitively, one might expect training on larger, more diverse image pools to improve generalization. However, our results demonstrate that even agents trained on pools of up to 100 images completely fail to transfer skills to novel images in the Hard OOD setting. Furthermore, performance on augmented training images often decreases as training pool size increases. This counter-intuitive result may suggest that agents trained on smaller, less diverse pools (which are often more sample efficient) learn representations more attuned to the specific structural invariances of the task, making them robust to simple perturbations but not fundamentally general. The rare non-zero success rates in Hard OOD likely stem from chance encounters with nearly-solved initial states rather than genuine generalization.
@@
These findings offer critical insights for developing future methods. The strong link between Easy OOD robustness and sample efficiency suggests that encouraging invariance to simple transformations is beneficial. However, the universal failure on Hard OOD, and the degradation of Easy OOD with increasing pool diversity, indicates that current methods do not learn sufficiently abstract or disentangled representations. Future research should focus on developing techniques that not only achieve good sample efficiency but also explicitly promote generalization to truly novel visual contexts. This might involve architectures that better separate visual representation learning from policy learning, incorporate stronger inductive biases for visual reasoning, or leverage self-supervised objectives that encourage learning of more fundamental visual features. Simply increasing the diversity of training images, as SPGym allows, is insufficient with current algorithms, highlighting the need for new approaches to bridge this generalization gap. Full OOD performance data for both Easy and Hard settings across all configurations can be found in Appendix C.2.
@@ 5.4 Is representation quality linked with performance?
To measure the quality of learned representations, we performed linear probing on frozen encoders from trained PPO and SAC agents, using a single-layer MLP to predict one-hot puzzle states. We find a statistically significant correlation between probe accuracy and sample efficiency (Pearson r=-0.81, p=1.1e-13), indicating that encoders capturing more task-relevant spatial information are strongly linked with faster learning. As image pool size increases, both probe accuracy and task performance systematically degrade, with standard SAC maintaining high accuracy (100% at pool size 1, 97.63% at pool size 5) mirroring its strong efficiency, while less efficient methods like SAC+VAE (78.21% at pool size 5) and SPR (dropping from 94.31% to 75.48% from pool size 5 to 10) show reduced probe performance. These consistent trends across algorithms demonstrate SPGym’s ability to identify learning procedures that develop representations better aligned with the task’s spatial reasoning needs. Full results are provided in Appendix C.1.2.
@@ 5.5 Does performance generalize across image sources?
To validate that our findings generalize beyond ImageNet, we evaluated agents on DiffusionDB (Wang et al., 2023), a dataset of procedurally generated images. As shown in Figure 8, performance scaling patterns on DiffusionDB closely mirror those observed on ImageNet across PPO, SAC, and DreamerV3. This consistency across fundamentally different image sources, real photographs versus synthetic generations, demonstrates that visual diversity rather than semantic content drives the representation learning challenge in SPGym. The similarity in degradation patterns as pool size increases indicates that our algorithmic insights reflect fundamental properties of the tested methods rather than dataset-specific artifacts. Procedurally generated images also offer practical advantages for future research: eliminating storage requirements through on-demand generation, enabling fine-grained control over visual similarity, and providing unlimited training diversity. For a more detailed analysis, see Appendix A.5.
@@ 5.6 How does puzzle size affect learning performance?
Another direction is increasing puzzle size. As shown in Table 3, the complexity increase from $3\times 3$ to $4\times 4$ grids significantly impacts learning. On the simpler $3\times 3$ puzzle, PPO solved the puzzle in 1.75M steps, while SAC and DreamerV3 were more efficient, requiring 0.33M and 0.42M steps, respectively. For the $4\times 4$ puzzle, PPO’s sample requirements surged to 24.46M steps, far exceeding the 10M step training budget, while SAC and DreamerV3 still solved the puzzle within budget at 8.14M and 2.26M steps, respectively. This demonstrates that while larger state spaces pose a major challenge by requiring more exploration and more complex visual representations, more sample-efficient algorithms can still scale to harder tasks.
@@ 6 Conclusion
We introduce the Sliding Puzzles Gym (SPGym), a novel benchmark designed to systematically evaluate representation learning in RL algorithms by isolating visual complexity from environment dynamics. Our analysis reveals three fundamental tensions in visual RL that challenge current methods. First, many sophisticated representation learning techniques struggle with SPGym’s unique combination of visual diversity and structured spatial reasoning, with auxiliary objectives often underperforming compared to simpler approaches like data augmentation. Second, we uncover a complex relationship between sample efficiency and generalization, where agents trained on smaller pools often learn representations more attuned to task-specific invariances, making them paradoxically more robust to simple perturbations. Most critically, we expose severe limitations in current end-to-end RL methods through their universal failure on completely novel visual contexts, achieving near-zero success rates despite strong in-distribution performance – revealing that learned representations rely primarily on memorization rather than genuine visual understanding. These findings highlight the need for fundamental algorithmic advances that effectively balance sample efficiency with true visual generalization.

### v5

@@ 4 Methods
In this section, we detail the evaluation protocol for various RL agents within SPGym, focusing on their ability to handle increasing visual diversity across training runs.
@@
We sample images from ImageNet-1k’s validation split (Russakovsky et al., 2015) to construct visual observations, resizing each image to $84\times 84$ pixels and normalizing values to $[0,1]$. To isolate visual diversity, we fix the puzzle size to $3\times 3$ and vary only the image pool size $p$. For each training run, we randomly sample $p$ distinct images to create a fixed image pool $\mathcal{I}$. At episode start, we randomly select one image from this pool to generate the puzzle observations. We cap the number of environment steps to 10M and limit episodes to 1,000 steps.
@@
For SAC, we begin with the standard implementation for discrete action spaces proposed by Christodoulou (2019). We then examine several representation learning variants: RAD (Laskin et al., 2020a), which employs data augmentation; CURL (Laskin et al., 2020b), which uses contrastive learning; SPR (Schwarzer et al., 2020), which incorporates self-supervised prediction; DBC (Zhang et al., 2021), which focuses on state metric learning; SAC-AE and SAC-VAE (Yarats et al., 2021b), which utilize reconstruction-based learning; and a simple baseline (SB) proposed by Tomar et al. (2023), which implements a simplified approach with reward and transition prediction.
@@
For SAC-based agents, we follow Yarats et al. (2021b) and Tomar et al. (2023) by blocking actor gradients through the encoder to prevent representation collapse, while allowing critic and auxiliary gradients. PPO and DreamerV3 maintain their original gradient flow patterns. While PPO and DreamerV3 worked robustly with default configurations, SAC-based agents required tuning of the temperature parameter $\alpha$ (Appendix A.4.1). We found a fixed value of $0.05$ to work best across all SAC variants, as the automatic tuning from Haarnoja et al. (2018b) proved ineffective here.
Where applicable, we preserve uniform hyperparameters while drawing algorithm-specific configurations from their respective source papers. The publicly available code repository [^2] Accessible at https://github.com/bryanoliveira/spgym-experiments extends both CleanRL (Huang et al., 2022) and the official DreamerV3 codebase. Additional methodological details are provided in Appendix A.
@@ 5 Results
5.1 Can SPGym distinguish agents on their approaches to representation learning?
Table 1: Million steps to reach 80% success rate across pool sizes during training. Lower is better. Best performing variant for each algorithm and pool size is highlighted in bold.
@@ 5.1 Can SPGym distinguish agents on their approaches to representation learning?
Table 1 demonstrates SPGym’s ability to differentiate agents based on their representation learning approaches across varying levels of visual diversity. Our primary goal in this analysis is to compare the sample efficiency with which different agents learn useful visual representations.
@@
To explore the potential benefits of leveraging pretrained encoders in this setup, we evaluate PPO with two types of pretrained encoders. It is important to acknowledge that direct comparisons of sample efficiency with encoders trained from scratch can be nuanced, as pretrained encoders have inherently been exposed to significantly more data during their pretraining phase. The in-distribution pretrained encoder (PT (ID)) for PPO, intended as a proxy to estimate an approximate upper bound on performance achievable with pretraining, significantly boosts sample efficiency across all tested pool sizes. The out-of-distribution pretrained encoder (PT (OOD)), designed to represent the average performance one might expect from a generally pretrained encoder, offers more modest gains which also decrease with larger pools. Notably, however, the rate at which performance degrades as the image pool size increases appears to be similar regardless of whether pretraining is used or not.
For SAC, data augmentation with RAD consistently improves efficiency, especially as visual diversity increases. In contrast, auxiliary methods like CURL, DBC, VAE, SPR, and SB require more samples than standard SAC, particularly with larger pools. The underperformance of these methods appears to stem from different factors. CURL may struggle because enforcing similarity between augmented observations could be ineffective when images from the pool lack shared structure. DBC might perform poorly when observations with similar visual features correspond to different dynamics and rewards. SPR, SB, and VAE may also be less effective because enforcing latent space smoothness could be unsuited to SPGym’s highly discontinuous observations, as suggested by the relatively better performance of the unconstrained AE variant.
@@
Finally, DreamerV3 demonstrates particularly strong performance, consistently outperforming both PPO and SAC variants across all pool sizes. The full DreamerV3 model shows remarkably stable performance across different pool sizes, significantly better than most other methods. The variant without decoder gradients shows reduced performance across all configurations, highlighting the importance of the world model’s discrete reconstruction objective. This suggests that DreamerV3’s approach of learning a predictive model of the environment provides a particularly effective foundation for handling visual diversity in SPGym.
@@
These results highlight SPGym’s value as a diagnostic tool, effectively distinguishing which approaches can handle its distinct combination of visual diversity and structured dynamics. The benchmark exposes important limitations in methods that have proven successful in other visual RL domains, suggesting that their underlying assumptions may not align with SPGym’s unique characteristics. For detailed performance analysis and proposed explanations for agent behavior, see Appendix C.2.1.
@@ 5.2 How does visual diversity affect agents?
These distinct degradation patterns suggest differences in handling increasing visual diversity, potentially due to agents memorizing features instead of learning generalizable representations, thus exhausting network capacity. PPO’s early failure indicates difficulty forming robust representations and rapid capacity exhaustion. SAC’s better performance on medium pools, possibly aided by its replay buffer and encoder training independent of actor gradients, suggests stronger representation learning initially. However, it also faces fundamental limitations, potentially as memorization becomes unsustainable. DreamerV3’s world model architecture appears to foster more compressed representations, leading to more graceful degradation, but it too struggles with very large pools.
Preliminary DreamerV3 experiments with extremely large image pools (10,000-20,000 images) show that agents fail to learn useful policies, as each observation becomes virtually unique. This suggests that simply increasing dataset scale is not enough if the RL signal becomes too sparse to effectively guide encoder training in the presence of high diversity. These methods appear to memorize rather than generalize, struggling when network capacity is overwhelmed by constantly novel inputs. This highlights fundamental limitations of current approaches when faced with diverse visuals where assumptions break and the learning signal is too weak to support meaningful learning.
@@ 5.3 Can agents generalize to unseen visual inputs?
Our analysis of generalization capabilities on PPO and SAC agents reveals a stark contrast between in-distribution and out-of-distribution test performance. We evaluate two levels of generalization difficulty: ‘Easy’ OOD, using augmented versions of training images, and ‘Hard’ OOD, using completely unseen pictures. Despite most agents achieving high success rates on their training distribution, they exhibit varying degrees of failure on both generalization challenges. Since meaningful generalization evaluation requires agents that first achieve competent in-distribution performance, we primarily focus our OOD analysis on smaller pool sizes (i.e., 1, 5, and 10) where most methods successfully learn the task. Full OOD performance data for both Easy and Hard settings across all configurations can be found in Appendix C.2.
Table 2: Success rate of PPO and SAC agents on Easy OOD across different training image pool sizes. Higher is better.
@@
For Easy OOD evaluation, we measure the success rate of PPO and SAC agents after applying each augmentation described in Appendix A.3, averaging results over 100 episodes per augmentation. Table 2 reports these results for training pool sizes 1, 5, and 10. As the diversity of the training pool increases, we observe a general decline in OOD performance. Some SAC variants show greater robustness to augmentations than standard SAC, especially when trained with a pool size of 1, but this advantage diminishes as the pool size grows. Overall, Easy OOD performance tends to decrease for all methods as training pool size increases. Notably, there is a strong correlation (Pearson r=0.81r=-0.81, $p=2.5\times 10^{-12}$) between Easy OOD success and sample efficiency across methods and pool sizes, suggesting that agents which scale well with increasing visual diversity also tend to generalize better to augmented inputs.
@@
For Hard OOD evaluation, we test all 5 seeds of PPO, SAC, and Dreamer agents on 100 episodes using images independently sampled from the ImageNet-1k validation split. Unsurprisingly, agents almost universally fail, achieving near 0% success rates across all methods and training configurations. This stark failure on completely unseen images is a crucial diagnostic finding, exposing fundamental limitations of current end-to-end RL methods when faced with visual inputs that differ substantially from their training distribution. This finding reinforces our interpretation that agents may be memorizing specific visual patterns instead of learning generalizable representations.
@@
Intuitively, one might expect training on larger, more diverse image pools to improve generalization. However, our results demonstrate that even agents trained on pools of up to 100 images completely fail to transfer skills to novel visuals in the Hard OOD setting. Furthermore, performance on augmented training images often decreases as training pool size increases. This perhaps counter-intuitive result may suggest that, in SPGym, agents trained on smaller, less diverse pools, learn representations more attuned to the specific structural invariances of the task, making them robust to simple perturbations but still not fundamentally general. The rare non-zero success rates in Hard OOD likely stem from chance encounters with nearly-solved initial states rather than genuine generalization.
@@
These findings reveal fundamental limitations in current visual RL methods and point toward specific research directions. The universal failure on Hard OOD, combined with degradation on Easy OOD as pool diversity increases, indicates that memorization-based learning is insufficient for visual generalization. Future research should focus on developing techniques that not only achieve good sample efficiency but also explicitly promote generalization to truly novel visual contexts. This might involve architectures that better separate visual representation learning from policy learning, incorporate stronger inductive biases for visual reasoning, or leverage self-supervised objectives that encourage learning of more fundamental visual features, along with regularization methods that explicitly discourage memorization. Simply increasing the diversity of training images, as SPGym allows, is insufficient with current algorithms, highlighting the need for new approaches.
@@ 5.4 Is representation quality linked with performance?
As a proxy for the quality of learned representations, we performed linear probing on frozen encoders from trained PPO and SAC agents, using a single-layer MLP to predict one-hot puzzle states. We find a statistically significant correlation between probe accuracy and sample efficiency (Pearson r=0.81r=-0.81, $p=1.1\times 10^{-13}$), indicating that encoders capturing more task-relevant spatial information are strongly linked with faster learning. As image pool size increases, both probe accuracy and task performance systematically degrade, with standard SAC maintaining high accuracy (100% at pool size 1, 97.63% at pool size 5) mirroring its strong efficiency, while less efficient methods like SAC+VAE (78.21% at pool size 5) and SPR (dropping from 94.31% to 75.48% from pool size 5 to 10) show reduced probe performance. These consistent trends across algorithms demonstrate SPGym’s ability to identify learning procedures that develop representations better aligned with the task’s spatial reasoning needs. Detailed probing results are provided in Appendix C.1.2.
@@ 5.5 Does performance generalize across image sources?
To validate that our findings generalize beyond ImageNet, we evaluated agents on DiffusionDB (Wang et al., 2023), a dataset of procedurally generated images. As shown in Figure 8 in Appendix A.5, performance scaling patterns on DiffusionDB closely mirror those observed on ImageNet across PPO, SAC, and DreamerV3. This consistency across fundamentally different image sources (real photographs versus synthetic generations) demonstrates that visual diversity rather than semantic content drives the representation learning challenge in SPGym. The similarity in degradation patterns as pool size increases indicates that our algorithmic insights reflect fundamental properties of the tested methods rather than dataset-specific artifacts. Procedurally generated images also offer practical advantages for future research: eliminating storage requirements through on-demand generation, enabling fine-grained control over visual similarity, and providing unlimited training diversity.
@@ 5.6 How does puzzle size affect learning performance?
Beyond visual diversity, we also explore the impact of increasing puzzle size. As shown in Table 3, the complexity increase from $3\times 3$ to $4\times 4$ grids significantly impacts learning. On the simpler $3\times 3$ puzzle, PPO solved the puzzle in 1.75M steps, while SAC and DreamerV3 were more efficient, requiring 0.33M and 0.42M steps, respectively. For the $4\times 4$ puzzle, PPO’s sample requirements surged to 24.46M steps, far exceeding the 10M step training budget, while SAC and DreamerV3 still solved the puzzle within budget at 8.14M and 2.26M steps, respectively. This demonstrates that while larger state spaces pose a major challenge by requiring more exploration and more complex visual representations, more sample-efficient algorithms can still scale to harder tasks.
@@ 6 Conclusion
We introduce the Sliding Puzzles Gym (SPGym), a novel benchmark designed to systematically evaluate representation learning in RL algorithms by isolating visual complexity controls from the environment dynamics. Our analysis reveals fundamental tensions challenging current visual RL methods. We show that sophisticated representation learning techniques struggle with SPGym’s combination of visual diversity and structured spatial reasoning requirements, with auxiliary objectives often underperforming simpler approaches like data augmentation. We also uncover a complex relationship between sample efficiency and generalization: agents trained on smaller pools learn task-specific invariances that paradoxically make them more robust to simple perturbations, with larger pools often degrading generalization performance. Most critically, we expose severe limitations in end-to-end RL methods through universal failure on novel visual contexts, achieving near-zero success despite strong in-distribution performance – revealing that learned representations rely on memorization rather than genuine visual understanding. These findings highlight the need for algorithmic advances that move beyond memorization toward true visual understanding.